# Sharp Capacity Scaling of Spectral Optimizers in Learning Associative Memory

### 27 Mar 2026 v1

@@ 4.3 Optimality of Muon
A natural follow-up question to Theorem 3 is: is the $d^{1+\frac{1}{2\alpha}}$ one-step recovery rate optimal for the linear associative memory task, or can it be improved by choosing a different estimator $h({\mathbf{G}}_{0})$ of $\mathbf{W}$}? For example, our analysis used $\lambda\sim d^{-\alpha}$ while the limit $\lambda\to 0$ recovers the exact polar map. Can choosing such a sharper resolution improve our results?
In this section, we give a negative answer to this question by providing a heuristic argument for the one-step optimality of our stabilized variant of Muon. We first show that the Bayes optimal estimator must be spectrally equivariant, that is for the SVD of the gradient ${\mathbf{G}}_{0}={\mathbf{U}}{\mathbf{S}}{\mathbf{V}}^{\top}$, it holds that $h({\mathbf{G}}_{0})={\mathbf{U}}h({\mathbf{S}}){\mathbf{V}}^{\top}$ and $h({\mathbf{S}})$ is diagonal.
Proposition 6.
@@
Intuitively, this means that smaller singular values are blown up by a larger multiplicative factor. Note that we do not require $h$ itself to be monotonic. For example, the classical cubic Newton–Schulz iteration $h(z)=\frac{3}{2}z-\frac{1}{2}z^{3}$ [3] satisfies this assumption, as well as its higher-order iterates on the interval of convergence (see right figure) — see Appendix C.4 for details.
We now show that any $h$ satisfying Assumption 7 cannot improve the signal $\gamma_{ii}$. As in Eq. (5), we compute the strength of the signal via a first-order approximation of the auxiliary map $\phi$, and the slope is given via the singular values of the leave-one-out gradient ${\mathbf{G}}_{-i}=\sum_{j\neq i}q_{i}u_{i}v_{i}^{\top}$ as
@@
where the inequality follows from $\frac{h(s_{k})+h(s_{\ell})}{s_{k}+s_{\ell}}\leq\frac{h(s_{k})}{s_{k}}+\frac{h(s_{\ell})}{s_{\ell}}$, $\frac{h(s_{k})-h(s_{\ell})}{s_{k}-s_{\ell}}\leq\min\{\frac{h(s_{k})}{s_{k}},\frac{h(s_{\ell})}{s_{\ell}}\}$ and $h^{\prime}(s_{k})\leq\frac{h(s_{k})}{s_{k}}$ under Assumption 7. That is, the signal strength is roughly determined by how much the average singular value is blown up by $h$. If $h$ is Lipschitz, this is uniformly bounded by $\lVert h\rVert_{\operatorname{Lip}}$ (thus Eq. (6) is tight for $h_{\lambda}$). However, even without Lipschitz control, this is fundamentally limited by the average scale of the singular spectrum as we prove below.
Lemma 8.
@@
As such, we must have $\phi^{\prime}(0)\lesssim d^{\alpha}$ regardless of the choice of $h$, therefore the signal $\gamma_{ii}=\phi(q_{i})$ is upper bounded (ignoring higher-order terms) as $\phi(0)+\phi^{\prime}(0)q_{i}\lesssim\frac{1}{\sqrt{d}}+q_{i}d^{\alpha}$. In contrast, the noise and interaction terms $\phi(0)$ and $\gamma_{ij}$ are generally of size $\widetilde{\Theta}(1/\sqrt{d})$. As a consequence, we indeed require $q_{i}\gg d^{-\alpha-\frac{1}{2}}$, equivalently $i\ll d^{1+\frac{1}{2\alpha}}$ in the population regime to ensure recovery, matching the rate obtained in Theorem 3.
For example, Lemma 8 implies that taking the resolution as $\lambda\ll d^{-\alpha-1}$ instead of $\lambda\sim d^{-\alpha}$ in Theorem 3 essentially gives the exact polar map $h(z)\approx\operatorname{sign}(z)$, as this scale will never be ‘seen’ by the singular values. Nonetheless, even in this regime the average blowup of the singular values is of order $d^{\alpha}$, and so we will not see any improvement from using the polar map. In fact, from this argument we expect roughly the same recovery rate, which is indeed what we observe in our experiments in Section 6.
5 Multiple Steps of Muon
We now turn our attention to the entire update trajectory of Muon. To study the macroscopic scaling behavior of this process, we will adopt a simplifying heuristic. At step $t$, suppose all items $i=1,\cdots,d_{t}$ have been recovered with $\hat{p}_{t}(i\mid i)\approx 1$. We presume all items $i>d_{t}$ have not been recovered at all, i.e. $\hat{p}_{t}(i\mid i)\ll 1$, and approximate the current gradient as
@@ 5 Multiple Steps of Muon
Theorem 9 (multi-step recovery of Muon).
@@
Thus for sufficiently large $B$, the recovery exponent $2-(1-\frac{1}{2\alpha})^{t}$ converges exponentially to the information-theoretic maximum with a fixed learning rate. In comparison, for multiple steps of SGD, we show a strictly suboptimal scaling law for any learning rate schedule.
Theorem 10 (multi-step recovery of SGD).
@@
We give a brief intuition for this phenomenon. In the non-orthogonal setting, each item $i>d_{t}$ to be recovered must compete with noise from the top individual unclassified items $j\sim d_{t}$ with large frequencies, as well as the aggregate fluctuation from the bulk of the unclassified items, which leads to the two thresholds in Theorem 10. Hence, Muon can be interpreted as effectively removing the first threshold by amplifying the bulk (but not top) singular directions. Once $d_{t}>d$, the gradient becomes relatively more isotropic and so the effect of orthogonalization is less pronounced; the second threshold becomes the limiting factor for both optimizers.
@@
The rates in Theorem 9 and Theorem 10 are given for a constant (or at most logarithmically diverging) horizon $T$, and hence can achieve $\Omega(d^{2-\epsilon})$ recovery rate for any $\epsilon>0$. As in Section 4.3, this scaling is likely near-optimal in this regime. However when $T$ grows further, we expect that the batch size dependency should more accurately scale as $(TB)^{1/\alpha}$, as this is the information-theoretic upper bound on the set of all observed items.
We also emphasize that the approximation in Eq. (7) is only a heuristic. In actuality, after $t$ steps of Muon, while items $i\geq\bar{d}_{t}:=d_{t}\operatorname{polylog}(d)$ do have near-uniform logits, for items $d_{t}\leq i\leq\bar{d}_{t}$ the predicted scores $\hat{p}_{t}(i\mid i)$ can range anywhere between $\frac{1}{N}$ and $1$. Moreover, these scores depend in a complicated manner on all embeddings $\{u_{j},v_{j}\}_{j\in[N]}$, so that the arguments used to prove Theorem 3 cannot be directly applied. Theoretically, this issue can be bypassed by assuming the existence of gaps in the power-law spectrum as in Li et al. [30], however we do not pursue this approach. Instead, we leave the precise end-to-end guarantee as a conjecture stated below, and empirically validate the predictions of Theorem 9 in Figure 4 (on the exact Muon iterates).
@@
The recovery and convergence rates of Theorem 9 also hold for the exact Muon iterates ${\mathbf{W}}_{t+1}={\mathbf{W}}_{t}+\eta h_{\lambda_{t}}({\mathbf{G}}_{t})$.
@@ 6 Experiments
(a) Gradient descent (population).
(b) Muon (population).
Figure 2: Capacity scaling after one population Muon and GD step. We set $N=100,000$ and vary $d,\alpha$. Each experiment is repeated 16 times. For each $\alpha$, we fit the dimension exponents of the mean capacity $d^{C_{\alpha}}$ (dashed lines), and then find the best fit of exponents $C_{\alpha}$ in the form of $C_{\alpha}=c_{1}+\frac{c_{2}}{\alpha}$ (solid lines). Observe that Muon achieves much higher storage than GD, and the exponents are consistent with Theorems 3 and 5.
@@ 6.1 Linear Associative Memory
Figure 2 shows the storage capacity scaling of the memory matrix $\mathbf{W}$} after a single population Muon or GD step. We fix the vocabulary size at N=100, [^0] N=100{,}000 and vary both the power law exponent $\alpha$ and the embedding dimension $d$. Muon (Figure 2(b)) indeed achieves a dramatically larger storage capacity than GD (Figure 2(a)). Moreover, the fitted scaling exponents (bottom right) agree with our theoretical predictions in the population limit: Muon stores $d^{1+\frac{1}{2\alpha}}$ items (Theorem 3), whereas GD stores only $d^{\frac{1}{2\alpha}}$ items (Theorem 5).
@@
In Figure 4, we examine the multi-step capacity scaling of population Muon to test the predictions of Theorem 9 (which assumes the deflation heuristic). We fix N=250, [^0] N=250{,}000, vary $d$ and $\alpha$, and run Muon with $\lambda_{t}=0$ on the population cross-entropy objective for $T$ steps using a fixed learning rate $\eta\asymp\sqrt{d}$. After each step, we measure the storage capacity and fit a power law to extract its scaling exponent in $d$. As shown in Figures 4(a), 4(b), 4(c), the capacity increases with the number of training steps. Moreover, Figure 4(d) shows that, after sufficiently many steps, the weight matrix approaches the optimal $\widetilde{\Theta}(d^{2})$ capacity [38]. Figure 4(e) compares the fitted capacity exponents for all $(T,\alpha)$ pairs against the predictions of Theorem 9. We find good overall agreement, with larger deviations at smaller $\alpha$ and larger $T$ where non-asymptotic effects are expected to be more pronounced. Overall, these results suggest that the heuristic approximation we introduced in Eq. (7) captures the scaling behavior of the training dynamics reasonably well.
@@
Figure 4: Capacity after $T$ Muon steps on the population cross-entropy loss. We set N=250, [^0] N=250{,}000, $\eta=2\sqrt{d}$. Figures 4(a), 4(b), 4(c) report the capacity at $T=2,3,4$, respectively (see Figure 2(b) for $T=1$); Figure 4(d) presents the capacity at large $T$: we run Muon for up to $500$ steps and early stop when the capacity improvement over $10$ steps drops below $0.5\%$. Figure 4(e) compares the fitted dimension exponents against predictions of Theorem 9; observe that the exponents agree except at small $\alpha$ and large $T$.
@@
Figure 5: Capacity scaling of multi-step Muon and GD. We set $N=100,000$, $\alpha=1.5$. (a) Population update: for GD we implement an increasing learning rate schedule (see Theorem 10) with $\eta_{1}=0.01\sqrt{d}$; for Muon we use a fixed step size $\eta=\sqrt{d}$. Observe that the benefit of Muon is most visible in the “early phase” of training (the initial plateau of GD in the first 3 steps is due to small $\eta_{1}$ chosen for numerical stability). (b) Capacity of minibatch Muon vs. total sample size $B\times T$; for each batch size $B$, we run minibatch Muon for $T=20$ steps with $\eta=\sqrt{d}$. Dashed red line indicates the information-theoretic rate $(BT)^{1/\alpha}$.
In Figure 5, we further compare the multi-step capacity scaling of GD and Muon with N=100, [^0] N=100{,}000 and $\alpha=1.5$. Figure 5(a) compares their performance in minimizing the population cross-entropy loss. Muon attains much higher capacity than GD in the first few steps. On the other hand, with the increasing learning rate schedule from Theorem 10, GD catches up later in training, and both methods eventually reach the optimal $d^{2}$ capacity (we however note that this increasing $\eta_{t}$ schedule for GD is numerically unstable, especially when $\alpha$ is large). These results suggest that Muon’s acceleration is most significant early in training, matching the predictions in Theorems 9 and 10.
Finally, Figure 5(b) shows minibatch Muon up to $T=20$ steps with batch sizes $B\in\{2^{11}\!,\dots,2^{16}\}$. For batch sizes below the critical threshold, the capacity stays close to the information-theoretic limit $(BT)^{1/\alpha}$, i.e., the number of items that can be observed after $T$ steps with batch size $B$. For larger batch sizes (darker curves), the capacity saturates and sample efficiency worsens.

### 28 Apr 2026 v2

@@ 5.1 Optimality of Muon
A natural follow-up question to Theorem 4.1 is: is the $d^{1+\frac{1}{2\alpha}}$ one-step recovery rate optimal among first-order methods for the linear associative memory task, or can it be improved by choosing a different estimator $h({\mathbf{G}}_{0})$ of $\mathbf{W}$}? For example, our analysis used $\lambda\sim d^{-\alpha}$ while the limit $\lambda\to 0$ recovers the exact polar map; can choosing a sharper resolution improve our results?
In this section, we give a negative answer to this question by providing a heuristic argument for the one-step optimality of our stabilized variant of Muon; this intuitively aligns with the fact that Muon already matches the recovery rate of Newton’s method (Theorem 4.4). We first show that any Bayes optimal gradient-based estimator must be spectrally equivariant, that is for the SVD of the gradient ${\mathbf{G}}_{0}={\mathbf{U}}{\mathbf{S}}{\mathbf{V}}^{\top}$, it holds that $h({\mathbf{G}}_{0})={\mathbf{U}}h({\mathbf{S}}){\mathbf{V}}^{\top}$ and $h({\mathbf{S}})$ is diagonal.
Proposition 5.1.
@@
Intuitively, this means that smaller singular values are blown up by a larger multiplicative factor. Note that we do not require $h$ itself to be monotonic. For example, the classical cubic Newton–Schulz iteration $h(z)=\frac{3}{2}z-\frac{1}{2}z^{3}$ [3] satisfies this assumption, as well as its higher-order iterates on the interval of convergence (see right figure) — see Appendix D.3 for details.
We now show that any $h$ satisfying Assumption 5.2 cannot improve the signal $\gamma_{ii}$. As in Eq. (8), we compute the strength of the signal via a first-order approximation of the auxiliary map $\phi$, and the slope is given via the singular values of the leave-one-out gradient ${\mathbf{G}}_{-i}=\sum_{j\neq i}q_{i}u_{i}v_{i}^{\top}$ as
@@
where the inequality follows from $\frac{h(s_{k})+h(s_{\ell})}{s_{k}+s_{\ell}}\leq\frac{h(s_{k})}{s_{k}}+\frac{h(s_{\ell})}{s_{\ell}}$, $\frac{h(s_{k})-h(s_{\ell})}{s_{k}-s_{\ell}}\leq\min\{\frac{h(s_{k})}{s_{k}},\frac{h(s_{\ell})}{s_{\ell}}\}$ and $h^{\prime}(s_{k})\leq\frac{h(s_{k})}{s_{k}}$ under Assumption 5.2. That is, the signal strength is roughly determined by how much the average singular value is blown up by $h$. If $h$ is Lipschitz, this is uniformly bounded by $\lVert h\rVert_{\operatorname{Lip}}$ (thus Eq. (9) is tight for $h_{\lambda}$). However, even without Lipschitz control, this is fundamentally limited by the average scale of the singular spectrum as we prove below.
Lemma 5.3.
@@
As such, we must have $\phi^{\prime}(0)\lesssim d^{\alpha}$ regardless of the choice of $h$, therefore the signal $\gamma_{ii}=\phi(q_{i})$ is upper bounded (ignoring higher-order terms) as $\phi(0)+\phi^{\prime}(0)q_{i}\lesssim\frac{1}{\sqrt{d}}+q_{i}d^{\alpha}$. In contrast, the noise and interaction terms $\phi(0)$ and $\gamma_{ij}$ are generally of size $\widetilde{\Theta}(1/\sqrt{d})$. As a consequence, we indeed require $q_{i}\gg d^{-\alpha-\frac{1}{2}}$, equivalently $i\ll d^{1+\frac{1}{2\alpha}}$ in the population regime to ensure recovery, matching the rate obtained in Theorem 4.1.
For example, Lemma 5.3 implies that taking the resolution as $\lambda\ll d^{-\alpha-1}$ instead of $\lambda\sim d^{-\alpha}$ in Theorem 4.1 essentially gives the exact polar map $h(z)\approx\operatorname{sign}(z)$, as this scale will never be ‘seen’ by the singular values. Nonetheless, even in this regime the average blowup of the singular values is of order $d^{\alpha}$, and so we will not see any improvement from using the polar map. In fact, from this argument we expect roughly the same recovery rate, which is indeed what we observe in our experiments in Section 6.
5.2 Multiple steps of Muon
We now turn our attention to the entire update trajectory of Muon and SGD. To study the macroscopic scaling behavior of these processes, we will adopt a simplifying heuristic. At step $t$, suppose all items $i=1,\cdots,d_{t}$ have been recovered with $\hat{p}_{t}(i\mid i)\approx 1$. We presume all items $i>d_{t}$ have not been recovered at all, i.e. $\hat{p}_{t}(i\mid i)\ll 1$, and approximate the current gradient as
@@ 5.2 Multiple steps of Muon
Theorem 5.4 (multi-step recovery of Muon).
@@
Thus for sufficiently large $B$, the recovery exponent $2-(1-\frac{1}{2\alpha})^{t}$ converges exponentially to the information-theoretic maximum $2$ with a fixed learning rate. In comparison, for multiple steps of SGD, we show a strictly suboptimal scaling law for any learning rate schedule.
Theorem 5.5 (multi-step recovery of SGD).
@@
We give a brief intuition for this phenomenon. In the non-orthogonal setting, each item $i>d_{t}$ to be recovered must compete with noise from the top individual unclassified items $j\sim d_{t}$ with large frequencies, as well as the aggregate fluctuation from the bulk of the unclassified items, which leads to the two thresholds in Theorem 5.5. Hence, Muon can be interpreted as effectively removing the first threshold by amplifying the bulk (but not top) singular directions. Once $d_{t}>d$, the gradient becomes relatively more isotropic and so the effect of orthogonalization is less pronounced; the second threshold becomes the limiting factor for both optimizers.
@@
The rates in Theorem 5.4 and Theorem 5.5 are given for a constant (or at most logarithmically diverging) horizon $T$, and hence can achieve $\Omega(d^{2-\epsilon})$ recovery rate for any $\epsilon>0$. As in Section 5.1, this scaling is likely near-optimal in this regime. However when $T$ grows further, we expect that the batch size dependency should more accurately scale as $(TB)^{1/\alpha}$, as this is the information-theoretic upper bound on the set of all observed items after $T$ steps.
We also emphasize that the approximation in Eq. (10) is heuristic. Under the exact Muon dynamics after $t$ steps, items with indices $i\geq\bar{d}_{t}:=d_{t}\operatorname{polylog}(d)$ have nearly uniform logits, but for items in the intermediate range $d_{t}\leq i\leq\bar{d}_{t}$, the predicted scores $\hat{p}_{t}(i\mid i)$ can take any value between $\frac{1}{N}$ and $1$. These scores also depend in a complicated way on all embeddings $\{u_{j},v_{j}\}_{j\in[N]}$, preventing a direct extension of the proof of Theorem 4.1. One way to bypass this is to assume gaps in the power-law spectrum as in Li et al. [31], but we do not take this route. Instead, we leave the precise end-to-end guarantee as a conjecture below and empirically validate predictions of Theorem 5.4 in Figure 4 (on the exact Muon iterates). From Theorem 4.4, we also conjecture that Muon continues to match Newton’s method throughout training, suggesting an intrinsic curvature-aware property.
@@
The recovery and convergence rates of Theorem 5.4 also hold for the exact Muon iterates ${\mathbf{W}}_{t+1}={\mathbf{W}}_{t}+\eta h_{\lambda_{t}}({\mathbf{G}}_{t})$. Moreover, Muon matches Newton’s method throughout training.
@@ 6.1 Linear associative memory
Figure 2 shows the storage capacity scaling of the memory matrix $\mathbf{W}$} after a single population Muon or GD step. We fix the vocabulary size at N=100, [^0] N=100{,}000 and vary both the power law exponent $\alpha$ and the embedding dimension $d$. Muon (Figure 2(b)) indeed achieves a dramatically larger storage capacity than GD (Figure 2(a)). Moreover, the fitted scaling exponents (bottom right) agree with our theoretical predictions in the population limit: Muon stores $d^{1+\frac{1}{2\alpha}}$ items (Theorem 4.1), whereas GD stores only $d^{\frac{1}{2\alpha}}$ items (Theorem 4.3).
@@
(a) Gradient descent (population).
(b) Muon (population).
Figure 2: Capacity scaling after one population Muon and GD step. We set $N=100,000$ and vary $d,\alpha$. Each experiment is repeated 16 times. For each $\alpha$, we fit the dimension exponents of the mean capacity $d^{C_{\alpha}}$ (dashed lines), and then find the best fit of exponents $C_{\alpha}$ in the form of $C_{\alpha}=c_{1}+\frac{c_{2}}{\alpha}$ (solid lines). Observe that Muon achieves much higher storage than GD, and the exponents are consistent with Theorems 4.1, 4.3.
@@
In Figure 4, we examine the multi-step capacity scaling of population Muon to test the predictions of Theorem 5.4 (which assumes the deflation heuristic). We fix N=250, [^0] N=250{,}000, vary $d$ and $\alpha$, and run Muon with $\lambda_{t}=0$ on the population cross-entropy objective for $T$ steps using a fixed learning rate $\eta\asymp\sqrt{d}$. After each step, we measure the storage capacity and fit a power law to extract its scaling exponent in $d$. As shown in Figures 4(a), 4(b), 4(c), the capacity increases with the number of training steps. Moreover, Figure 4(d) shows that, after sufficiently many steps, the weight matrix approaches the optimal $\widetilde{\Theta}(d^{2})$ capacity [39]. Figure 4(e) compares the fitted capacity exponents for all $(T,\alpha)$ pairs against the predictions of Theorem 5.4. We find good overall agreement, with larger deviations at smaller $\alpha$ and larger $T$ where non-asymptotic effects are expected to be more pronounced. Overall, these results suggest that the heuristic approximation we introduced in Eq. (10) captures the scaling behavior of the training dynamics reasonably well.
@@
Figure 4: Capacity after $T$ Muon steps on the population cross-entropy loss. We set N=250, [^0] N=250{,}000, $\eta=2\sqrt{d}$. Figures 4(a), 4(b), 4(c) report the capacity at $T=2,3,4$, respectively (see Figure 2(b) for $T=1$); Figure 4(d) presents the capacity at large $T$: we run Muon for up to $500$ steps and early stop when the capacity improvement over $10$ steps drops below $0.5\%$. Figure 4(e) compares the fitted dimension exponents against predictions of Theorem 5.4; observe that the exponents agree except at small $\alpha$ and large $T$.
@@
Figure 5: Capacity scaling of multi-step Muon and GD. We set $N=100,000$, $\alpha=1.5$. (a) Population update: for GD we implement an increasing learning rate schedule (see Theorem 5.5) with $\eta_{1}=0.01\sqrt{d}$; for Muon we use a fixed step size $\eta=\sqrt{d}$. Observe that the benefit of Muon is most visible in the “early phase” of training (the initial plateau of GD in the first 3 steps is due to small $\eta_{1}$ chosen for numerical stability). (b) Capacity of minibatch Muon vs. total sample size $B\times T$; for each batch size $B$, we run minibatch Muon for $T=20$ steps with $\eta=\sqrt{d}$. Dashed red line indicates the information-theoretic rate $(BT)^{1/\alpha}$.
In Figure 5, we further compare the multi-step capacity scaling of GD and Muon with N=100, [^0] N=100{,}000 and $\alpha=1.5$. Figure 5(a) compares their performance in minimizing the population cross-entropy loss. Muon attains much higher capacity than GD in the first few steps. On the other hand, with the increasing learning rate schedule from Theorem 5.5, GD catches up later in training, and both methods eventually reach the optimal $d^{2}$ capacity (we however note that this increasing $\eta_{t}$ schedule for GD is numerically unstable, especially when $\alpha$ is large). These results suggest Muon’s acceleration is most significant early in training, matching the predictions in Theorems 5.4 and 5.5.
Figure 5(b) shows minibatch Muon up to $T=20$ steps with batch sizes $B\in\{2^{11}\!,\dots,2^{16}\}$. For batch sizes below the critical threshold, the capacity stays close to the information-theoretic limit $(BT)^{1/\alpha}$, i.e., the number of items that can be observed after $T$ steps with batch size $B$. For larger batch sizes (darker curves), the capacity saturates and sample efficiency worsens.
(a) Muon vs. Newton ($\kappa=0$).
(b) Muon vs. Newton ($\kappa=1$).
(c) Muon vs. Newton ($\kappa=1.5$).
Figure 6: Capacity scaling after one (population) Muon and Newton step in the anisotropic setting: we choose $u_{i}\sim\mathcal{N}(0,\frac{1}{d}{\mathbf{I}}_{d})$, $v_{i}\sim\mathcal{N}(0,\boldsymbol{\Xi}_{v})$, where $\boldsymbol{\Xi}_{v}$ is a trace-normalized diagonal matrix with $\lambda_{i}(\boldsymbol{\Xi}_{v})\asymp i^{-\kappa}$, $\kappa\geq 0$. We set $N=100,000$ and vary $d,\alpha$. For Newton’s method we add a ridge regularization $\lambda=10^{-8}$ for numerical stability when the preconditioner is rank-deficient. Observe that when $\kappa=0$ (isotropic, Figure 6(a)), Muon and Newton both achieve $d^{1+\frac{1}{2\alpha}}$ capacity, but as $\kappa$ increases (Figures 6(b), 6(c)), the performance of Muon worsens while the Newton update remains invariant.
Comparison with Newton’s method.
Figure 6 compares the storage efficiency of Muon and Newton’s update in the population setting. As argued in Section 4.2, Muon should match Newton’s method when the embedding and unembedding vectors are isotropic. To vary the anisotropy of the associative memory model, we consider $u_{i}\sim\mathcal{N}(0,\tfrac{1}{d}{\mathbf{I}}_{d}),v_{i}\sim\mathcal{N}(0,\boldsymbol{\Xi}_{v}),$ where the covariance matrix $\boldsymbol{\Xi}_{v}$ has eigenvalues $\lambda_{i}(\boldsymbol{\Xi}_{v})\asymp i^{-\kappa}$, $\kappa\geq 0$. We observe that in the isotropic case ($\kappa=0$), Muon matches the storage capacity of one Newton step; but as the data become more anisotropic, the gap between the two methods grows, and only Newton’s method retains the $d^{1+\frac{1}{2\alpha}}$ capacity.